I'm confused about negative numbers when you add or subtract them. The problem I am stuck on is five/sixths subtract negative two/thirds? Please help. Kathryn
Answer by checkley71(8403)

5/6-(-2/3) A MINUS TIMES A MINUS EQUALS A PLUS SO:
5/6+2/3 NOW FIND A COMMON DENOMINATOR AND ADD
5/6+4/6=9/6=1 3/6 OR 1 1/2 OR 1.5